The annual cost in dollars for removing p% of the toxic chemicals from a town’s water supply is given by the formula C(p) 500,000/100-p

a) Use the accompanying graph to estimate the cost for removing 90% and 95% of the toxic chemicals.
b) Use the formula to find C(99.5) and C(99.9).
c) What happens to the cost as the percentage of pollutants removed approaches 100%?

Answer by checkley77(12844)

C(p) 500,000/100-p
c(90)=500,000/(100-90)
c(90)=500,000/10
c(95)=50,000 cost for 90% removal.
c(95)=500,000/(100-95)
c(95)=500,000/5
c(95)=100,000 cosy of removal of 95&
c(99.5)=500,000/(100-99.5)
c(99.5)=500,000/.5
c(99.5)=1,000,000 cost of 99.5% removal
c(99.9)500,000/(100-99.9)
c(99.8)=500,000/.1
c(99.9)=5,000,000 cost of 999% removal.
As the removal approaches 100% the cost approached infinity.